OpenStudy (dallaslifebaby01):

i just need the answers for these questions 1. What is the equation of the line that has a slope of 4 and a y-intercept of -7? 2.What is an equation of the line that passes through (0, 8) and (4, 0)?

OpenStudy (astrophysics):

That's not how it works here...

OpenStudy (dallaslifebaby01):

I'm not using them for work

OpenStudy (astrophysics):

We don't know that also, you should be able to do the second one, as you've asked several questions similar to it, \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] m is the slope, and your \[(x_1, y_1)~~~ (x_2,y_2)\] are given respectively. For the first one use y = mx+b where m is slope and b is the y - intercept, this is the equation of a line.

OpenStudy (batman19991):

1) y=mx+b m=slope, so m=4 then plug in the x and y to find b, like so: -2=4(0)+b b=-2 SO y=4x-2

OpenStudy (batman19991):

i think idk

OpenStudy (astrophysics):

That's wrong, but you should let @DallasLifeBaby01 try it out on their own

OpenStudy (batman19991):

i never said i was good at math

OpenStudy (dallaslifebaby01):

lol thats ok

OpenStudy (dallaslifebaby01):

I'm working it out

OpenStudy (batman19991):

haha ok